A worker in a radiation lab recieves a whole-body radiation dose of 21 mrad. Her mass is 75 kg. The radiation delivered by alpha particles for which the RBE is 12.

1) What was the total energy absorbed by her body?

Eabsorbed =

2) What was her equivalent absorbed dose in mREM?

Doseeq=

3) As a normal employee, your allowable equivalent dose for an entire year is around 500 mREM. What percentage of of her annual dose did this employee receive?

4) What if the employee's body had absorbed the same amount of energy from this radiation, but the radiation has been in the form of electrons, β particles, which had an RBE of 13? What was her equivalent absorbed dose in mREM?

Doseeq=
Answer #1

(1) The total energy absorbed is equal to, = E =(Absorbed does)x(mass) = (21)*(75) = (21x10- rad)(75) = (21x10- J/kg)(75) =1.